# Repoformer: Selective Retrieval for Repository-Level Code Completion

### v1

@@ Abstract
Recent advances in retrieval-augmented generation (RAG) have initiated a new era in repository-level code completion. However, the invariable use of retrieval in existing methods exposes issues in both efficiency and robustness, with a large proportion of the retrieved contexts proving unhelpful or harmful to code language models (code LMs). To tackle the challenges, this paper proposes a selective RAG framework where retrieval is avoided when unnecessary. To power this framework, we design a self-supervised learning approach that enables a code LM to accurately self-evaluate whether retrieval can improve its output quality and robustly leverage the potentially noisy retrieved contexts. Using this LM as both the selective retrieval policy and the generation model, our framework consistently outperforms the state-of-the-art prompting with an invariable retrieval approach on diverse benchmarks including RepoEval, CrossCodeEval, and a new benchmark. Meanwhile, our selective retrieval strategy results in strong efficiency improvements by as much as 70% inference speedup without harming the performance. We demonstrate that our framework effectively accommodates different generation models, retrievers, and programming languages. These advancements position our framework as an important step towards more accurate and efficient repository-level code completion.
@@ 1 Introduction
Figure 1: (a) An overview of the proposed selective RAG framework. Given the current file context, a policy first assesses whether retrieval is required and triggers the retriever selectively. Then, the code LM makes a hypothesis with the optional retrieved context. With Repoformer, the two stages are streamlined via self-assessment. (b) With Repoformer acting as both the selective retrieval policy and the generator LM, our framework achieves better accuracy and better latency than performing retrieval invariably.
Automatic code completion has attracted long-lasting research efforts due to its high practical value in improving programmer productivity (Ye & Fischer, 2002; Hill & Rideout, 2004; Hellendoorn & Devanbu, 2017). One particularly challenging scenario is repository-level code completion, where a system is required to complete lines, API invocations, or functions in a file from user repositories. For this task, language models for code (code LMs) have emerged as a promising solution due to their ability to leverage the current file’s context to generate coherent code of flexible granularity (Tu et al., 2014; Svyatkovskiy et al., 2020; Chen et al., 2021). However, their performance is limited by the lack of repository-specific knowledge such as user-defined APIs and inter-module dependencies that require information beyond the current file to fully comprehend (Zan et al., 2022; Zhang et al., 2023; Ding et al., 2023). To bridge this gap, recent approaches adopt the retrieval-augmented generation (RAG) paradigm: cross-file contexts including code snippets or documentations are retrieved from the same repository and provided to code LMs as augmentations to the current file. Under this paradigm, existing works have advanced the retrieval mechanism for prompting black-box code LMs (Lu et al., 2022; Shrivastava et al., 2023b; Zhang et al., 2023) as well as adapted the LM to better leverage structured retrieved contexts such as classes, functions, or APIs (Ding et al., 2022; Zan et al., 2022).
@@
Our findings suggest that the answer is predominantly negative, primarily for two reasons. First, the benefit from retrieval is often sparse. On various code completion tasks, we discover that up to 80% of retrievals made by a standard RAG method do not enhance the performance of strong code LMs such as StarCoder (Li et al., 2023b), with many degrading performances by introducing irrelevant information (Section 5.1). Second, always retrieving introduces notable inefficiencies. The size of retrieval index grows linearly with the number of lines in the repository. For moderately sized repositories, sparse retrieval is already as time consuming as completion with a 3B LM (Section 5.3 and 6). This inefficiency is more pronounced with dense retrieval, enterprise-scale repositories, and iterative retrieval methods (Zhang et al., 2023). Hence, there is a critical need to reevaluate the“invariable retrieval” assumption. Such reevaluation is not only pivotal for enhancing code completion but also offers valuable insights for a broad range of similar applications that use domain-specific structured retrieval.
To tackle these challenges, we propose a novel repository-level code completion framework underpinned by a selective retrieval mechanism: the system uses a policy to proactively abstain from unnecessary or potentially detrimental retrievals (Figure 1 (a)). At the core of our framework is Repoformer, an intelligent code LM fine-tuned for robust code completion with self-triggered retrieval augmentation. Repoformer reflects three core principles:
@@
    Performance-oriented self-evaluation. After observing the current file, Repoformer explicitly expresses the likelihood that its prediction quality would be improved by additional retrieval. Our training strategy enables the model to combine two factors for abstention: the LM already knowing the answer without retrieval (Kadavath et al., 2022) and the code completion question not depending on cross-file information to answer and thus retrieval is likely uninformative.
@@
    Generalizability. The aforementioned two abilities must generalize to any completion granularity, language, and retriever. In addition, as a selective retrieval policy, Repoformer should be able to function as a plug-and-play policy for other black-box code LMs.
Correspondingly, we design a simple self-supervised RAG simulation procedure to create the training data. Diverse code chunks and function bodies are sampled and paired with the retrieved cross-file contexts. Then, the ground-truth label for selective retrieval can be obtained by contrasting the quality of a code LM’s outputs with and without retrieval augmentation. With this dataset, a multi-task objective is designed to jointly teach the code LM the ability to self-evaluate and robustly completing the code with optional retrieval augmentation (Section 3.3).
We perform comprehensive evaluations on a variety of repository-level code completion tasks from RepoEval (Zhang et al., 2023), CrossCodeEval (Ding et al., 2023), and a newly created large-scale benchmark. Results show that Repoformer achieves strong performance, outperforming invariable retrieval with the same-sized StarCoderBase by more than 3 absolute points for edit similarity across multiple tasks. The 3B Repoformer even performs on par with invariable retrieval using the 16B StarCoder (Section 5.2). Furthermore, our framework allows for up to 70% inference speedup without harming accuracy. We also establish that Repoformer can accelerate RAG with larger black-box code LMs as a plug-and-play selective RAG policy, improving the performance while reducing the latency of line and API completion to 75% (Section 5.3). Finally, we analyze Repoformer’s threshold sensitivity, the precision and calibration of its abstention decisions, its robustness to retrieved contexts, and its generalization to other languages or retrievers (Section 6). Our code, model, and the new evaluation benchmark will be released at https://repoformer.github.io/.
@@ 2 Related Work
Accurately completing the code in repositories has been a challenging research problem due to cross-file dependency patterns caused by modular design (Parnas, 1972; Tu et al., 2014). Early works propose application-specific training methods for n-gram LMs (Tu et al., 2014), RNNs (Hellendoorn & Devanbu, 2017; Wang et al., 2021), and Transformers (Svyatkovskiy et al., 2020) to leverage structured knowledge beyond current file’s context. Recent studies investigate fine-tuning powerful pre-trained code LMs (Chen et al., 2021; Nijkamp et al., 2022b; Li et al., 2023b) to better leverage retrieved knowledge provided in context such as code and documentation snippets (Zan et al., 2022; Ding et al., 2022; Shrivastava et al., 2023a). Concurrently, other studies show that black-box code LMs can already take advantage of in-context knowledge, depending on how well the knowledge is retrieved and formatted (Lu et al., 2022; Zhou et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023). This approach does not require one to train the LM and thus promises better generalization. Orthogonal to these studies, this paper identifies and addresses the robustness and efficiency issues caused by invariably performing the retrieval augmentation. Our solution takes the form of selective retrieval augmentation through self-assessment.
@@
This paper is consistent with the recent trend of making the RAG paradigm active and adaptive. A core question is finding an effective policy to decide when to retrieve. He et al. (2021) propose to learn to adjust the importance weight of retrieval based on language modeling performance. Li et al. (2023a) and Jiang et al. (2023) suggest that retrieval should be performed only when LMs have a high predictive uncertainty. Mallen et al. (2023) discover that retrieval can be avoided for popular facts. Concurrent to this work, two new studies approach adaptive RAG from a learning perspective. SKR (Wang et al., 2023) collects instances where retrieval is not helpful for black-box LMs and proposes several methods to predict these instances. Self-RAG (Asai et al., 2024) utilizes GPT-4 (OpenAI, 2023) as a knowledge engine to distill a smaller LM to evaluate whether answering a question can be benefited from retrieval. In comparison, this paper highlights the importance of understanding whether an LM knows the answer (Kadavath et al., 2022) in forming the retrieval policy. We introduce a simple yet effective scheme to fine-tune a code LM for faithful self-evaluation without extra modules (SKR), knowledge store (SKR), or labels generated by an oracle LM (Self-RAG). We show that our approach leads to no performance harms (Section 5.2), substantial speedup (Section 5.3), and a good abstention accuracy (Section 6).
@@ 3.1 Background
  • A code LM $\mathcal{M}$ that leverages $X_{l}$, $X_{r}$, and $CC$ to output $\hat{Y}$. The inclusion of $X_{r}$ and $CC$ is optional. It is worth noting that for generation, we always directly provide $X_{r}$ in the prompt in addition to $X_{l}$ (Shrivastava et al., 2023b; Pei et al., 2023). We provide empirical support for this design in Appendix B.
Full documentation of the RAG stages and their hyperparameters is provided in Appendix A for further reference.
@@ 3.2 Self-selective RAG for Code Completion
Central to our framework is the idea of selective RAG, where the system decides whether the LM’s generation could benefit from retrieved contexts and abstains from retrieval augmentation when it is deemed unnecessary (Figure 1 (a)).
For this selective decision, two traditional heuristics are relevant: (1) performing a trial retrieval and only augmenting the high-relevance contexts or (2) performing a trial generation and conducting RAG only when the model’s uncertainty is high. We find that these two strategies are informative to some extent: in line completion and API completion from RepoEval, both types of heuristics can achieve the performance of invariable retrieval with only 50% retrieval budget. However, they incur a high latency cost and do not generalize well to all tasks (Appendix C).
Instead, our framework adopts a self-selective RAG formulation. After observing $X_{l}$ and $X_{r}$, the LM directly self-triggers cross-file retrieval by generating a special token <cc> or abstains from retrieval via an empty token $\phi$ [^1] In practice, instead of greedily decoding <cc>, we check whether its probability exceeds a certain threshold.. This crucial formulation enables us to train a strong classifier that considers both the question’s characteristics (i.e., whether it requires cross-file information or not) and the LM’s self-knowledge (i.e., whether the LM can already correctly answer without retrieval, as explored by Kadavath et al. (2022)). After this optional retrieval step, the LM proceeds with the code completion with $X_{l}$, $X_{r}$, combined with $CC$ if retrieval is triggered.
As shown in Figure 2, the inference procedure of self-selective RAG is conveniently modeled as an extension to fill-in-the-middle (Bavarian et al., 2022). One advantage of this design is the flexibility. The LM possesses the ability for RAG and fill-in-the-middle, and can seamlessly self-switch between the two when encountering different questions. Users can also easily adjust the ratio between the two through the retrieval threshold. Another advantage is its efficiency. The selective decision overhead is only a single forward pass, a significant save compared to trial generation (such as Jiang et al. (2023)) or trial retrieval. When the LM abstains from retrieval, it can directly proceed with generation and the retrieval overhead is completely avoided.
Figure 2: A comparison between fill-in-the-middle and self-selective RAG. We mark the end of the current file with a new token <eof>, which triggers the LM’s self-evaluation. $\rightarrow$ denotes the invocation of the LM. LM-generated parts are colored in red. fim_p, fim_s, and fim_m refer to the special tokens for fill-in-the-middle: fim_prefix, fim_suffix, and fim_middle. These tokens are already learned during the pre-training.
@@ 3.3 Self-supervised Multi-task Learning
    Retrieve $CC$ using the current file, with or without $Y$.
@@
The detailed algorithm is presented in Appendix D. After running the algorithm, we obtain the fine-tuning instances, each in the form $(X_{l},X_{r},\ Y,\ CC,\ label)$.
@@ 4.1 Repoformer Implementation Details
We sample 18k Python repositories from the Stack [^2] We have also trained a version of Repoformer on a multilingual dataset. The results are reported in Section E.2. (Kocetkov et al., 2022) that have (1) at least three imports per file, (2) at least two local imports per file, and (3) at least five Python files. These criteria ensure the existence of local dependencies where RAG could be meaningful. We use $\mathcal{M}$ = StarCoderBase-1B and $T$ = 00 to label 240k chunk and 120k function completion instances. We reserve 500 repositories for validation and use the rest for training.
@@
We fine-tune the 1B and 3B variants of StarCoderBase with $\lambda=1.0$, maximum sequence length 2048, learning rate 2e-5, batch size 512, 50 warmup steps, and a linear learning rate decay. The models are trained for 2 epochs, which takes 8 hours for the 1B model and 12 hours for the 3B model with 8 Nvidia A100 GPUs (40G memory). We will call our models Repoformer-1B/3B.

### v2

@@ Abstract
Recent advances in retrieval-augmented generation (RAG) have initiated a new era in repository-level code completion. However, the invariable use of retrieval in existing methods exposes issues in both efficiency and robustness, with a large proportion of the retrieved contexts proving unhelpful or harmful to code language models (code LMs). In this paper, we propose a selective RAG framework to avoid retrieval when unnecessary. To power this framework, we design a self-supervised learning approach to enable a code LM to accurately self-evaluate whether retrieval can improve its output quality and robustly leverage the potentially noisy retrieved contexts. Using this LM as both the selective RAG policy and the generation model, our framework achieves state-of-the-art repository-level code completion performance on diverse benchmarks including RepoEval, CrossCodeEval, and CrossCodeLongEval, a new long-form code completion benchmark. Meanwhile, our analyses show that selectively retrieving brings as much as 70% inference speedup in the online serving setting without harming the performance. We further demonstrate that our framework is able to accommodate different generation models, retrievers, and programming languages. These advancements position our framework as an important step towards more accurate and efficient repository-level code completion.
@@ 1 Introduction
Figure 1: An overview of the proposed selective RAG framework. Given the current file context, the system first assesses whether retrieval is required and triggers the retriever if the question can likely be benefited from retrieval (right), abstaining from retrieval otherwise (left). Then, the code LM generates with optional retrieved contexts. With Repoformer, the two stages are streamlined via self-assessment.
Automatic code completion has attracted long-lasting research efforts due to its high practical value in improving programmer productivity (Ye & Fischer, 2002; Hill & Rideout, 2004; Hellendoorn & Devanbu, 2017). One particularly challenging scenario is repository-level code completion, where a system is required to complete lines, API invocations, or functions in a file from user repositories. For this task, language models for code (code LMs) have emerged as a promising solution due to their ability to leverage the context of the current file to generate coherent code of flexible granularity (Tu et al., 2014; Svyatkovskiy et al., 2020; Chen et al., 2021). However, these approaches fail to capture the holistic repository knowledge spanning beyond the current file, such as user-defined APIs and inter-module dependencies (Zan et al., 2022; Zhang et al., 2023; Ding et al., 2023). Recently, the retrieval-augmented generation (RAG) paradigm was proposed to bridge the gap: cross-file contexts such as relevant code snippets or documentations are retrieved and provided to code LMs as augmentations to the current file. This approach has shown strong empirical performance and was further advanced by recent literature through designing better retrieval mechanisms for prompting black-box code LMs (Lu et al., 2022; Shrivastava et al., 2023b; Zhang et al., 2023) and adapting the LM to better leverage structured retrieved contexts such as classes, functions, or APIs (Ding et al., 2024; Zan et al., 2022).
@@
Our findings suggest that the answer is predominantly negative. First, in various code completion tasks, we discover that up to 80% of the retrievals performed by a standard RAG method do not enhance the performance of common code LMs such as CodeGen (Nijkamp et al., 2023b) and StarCoder (Li et al., 2023b), and many degrade the performance by introducing irrelevant information (Section 5.1). Second, always retrieving introduces notable inefficiencies. For moderately sized repositories, sparse retrieval is already as time consuming as code completion with a 3B code LM (Section 5.3 and Section 6). This inefficiency is more pronounced with dense retrieval, enterprise-scale repositories, and iterative RAG methods such as Zhang et al. (2023).
In this paper, we challenge the assumption of always retrieving by proposing a novel repository-level code completion framework underpinned by a selective retrieval mechanism: the system proactively abstains from performing unnecessary or potentially detrimental retrievals (Figure 1 (a)). At the core of our framework is Repoformer, an intelligent code LM fine-tuned for robust code completion with self-triggered retrieval augmentation. Repoformer reflects three core principles:
@@
    Performance-oriented self-evaluation. After observing the current file, Repoformer explicitly expresses the likelihood that its prediction quality could be improved by cross-file retrieval. Our training strategy enables the model to combine two factors in this decision: the code LM already knowing the answer without retrieval (Kadavath et al., 2022) and the code completion question not depending on cross-file information and thus retrieval is likely uninformative.
@@
    Generalizability. The aforementioned two abilities must generalize to any completion granularity, programming language, and retriever choice. In addition, Repoformer should be able to function as a plug-and-play selective retrieval policy when other models are employed as the generation model.
We posit that these abilities can be faithfully obtained by learning from simulations of RAG. Specifically, we leverage a large number of permissively licensed repositories, sample diverse blanks to complete, and pair them with the retrieved repository-level cross-file contexts. Then, for a given code LM, the ground-truth label for selective retrieval is obtained by contrasting the quality of its outputs with and without retrieval augmentation. With this dataset, we design a self-supervised objective to jointly train code LMs to accurately self-evaluate the need for retrieval and robustly complete the code with the optional retrieval augmentation (Section 3.3).
We perform comprehensive evaluations on a range of repository-level code completion tasks from RepoEval (Zhang et al., 2023), CrossCodeEval (Ding et al., 2023), and CrossCodeLongEval a new large-scale benchmark focusing on code chunk and function completion. Results show that Repoformer achieves strong performance, outperforming always retrieving with the same-sized StarCoderBase by more than 3 absolute points for edit similarity across multiple tasks. The 3B Repoformer performs on par with always retrieving using the 16B StarCoder, and the 16B Repoformer achieves state-of-the-art performance across all the tasks (Section 5.2). Furthermore, our framework allows for up to 70% inference speedup without harming accuracy. We also establish that Repoformer can accelerate RAG with larger black-box LMs as a plug-and-play selective RAG policy, improving the performance while reducing the latency of line and API completion to 75% (Section 5.3).
Finally, in Section 6, we provide comprehensive analyses on Repoformer’s generalization ability. We show that Repoformer makes precise retrieval abstention decisions, is robust to retrieved contexts, and performs well when tested in other languages or with other retrievers. To facilitate future research on repository-level code completion, we will release our implementation and the CrossCodeLongEval benchmark at https://repoformer.github.io/.
@@ 2 Related Work
Accurately completing the code in repositories has been a challenging research problem due to cross-file dependency patterns caused by modular design (Parnas, 1972; Tu et al., 2014). Early works propose application-specific training methods for n-gram LMs (Tu et al., 2014), RNNs (Hellendoorn & Devanbu, 2017; Wang et al., 2021), and Transformers (Svyatkovskiy et al., 2020) to leverage structured knowledge beyond current file’s context. Recent studies investigate fine-tuning powerful pre-trained code LMs (Chen et al., 2021; Nijkamp et al., 2023b; Li et al., 2023b) to better leverage retrieved knowledge provided in context such as code and documentation snippets (Zan et al., 2022; Ding et al., 2024; Shrivastava et al., 2023a). Concurrently, other studies show that black-box code LMs can already take advantage of in-context knowledge, depending on how well the knowledge is retrieved and formatted (Lu et al., 2022; Zhou et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023). This approach does not require one to train the LM and thus promises better generalization. Orthogonal to these studies, this paper identifies and addresses the robustness and efficiency issues caused by invariably performing the retrieval augmentation. Our solution takes the form of selective retrieval augmentation through self-assessment.
@@
This paper is consistent with the recent trend of making the RAG paradigm active and adaptive. A core question is finding an effective policy to decide when to retrieve. He et al. (2021) propose to learn to adjust the importance weight of retrieval based on language modeling performance. Drozdov et al. (2022) proposes to upweight the retrieved information when the retrieval has high quality. Li et al. (2023a) and Jiang et al. (2023) suggest that retrieval should be performed only when LMs have a high predictive uncertainty. Mallen et al. (2023) discover that retrieval can be avoided for popular facts. Concurrent to this work, two new studies approach adaptive RAG from a learning perspective. SKR (Wang et al., 2023) collects instances where retrieval is not helpful for black-box LMs and proposes several methods to predict these instances. Self-RAG (Asai et al., 2024) utilizes GPT-4 (OpenAI, 2023) as a knowledge engine to distill a smaller LM to evaluate whether answering a question can be benefited from retrieval. In comparison, this paper highlights the importance of understanding whether an LM knows the answer (Kadavath et al., 2022) in forming the retrieval policy. We introduce a simple yet effective scheme to fine-tune a code LM for faithful self-evaluation without extra modules (SKR), knowledge store (SKR), or labels generated by an oracle LM (Self-RAG). We show that our approach leads to no performance harms (Section 5.2), substantial speedup (Section 5.3), and a high decision accuracy (Section 6).
@@ 3.1 Background
  • A code LM $\mathcal{M}$ that leverages $X_{l}$, $X_{r}$, and $CC$ to output $\hat{Y}$. The inclusion of $X_{r}$ and $CC$ is optional. In this paper, we always directly provide $X_{r}$ in the prompt in addition to $X_{l}$ (Shrivastava et al., 2023b; Pei et al., 2023). We provide empirical support for this design in Appendix B.
Full documentation of the RAG stages and their hyperparameters are provided in Appendix A for further reference.
@@ 3.2 Self-selective RAG for Code Completion
Central to our framework is the idea of selective RAG, where the system decides whether the LM’s generation could benefit from retrieved contexts and abstains from retrieval augmentation when it is deemed unnecessary (Figure 1).
For this selective decision, two traditional heuristics are relevant: (1) performing a trial retrieval and only augmenting the high-relevance contexts (e.g., Drozdov et al. (2022)) or (2) performing a trial generation and conducting RAG only when the model’s uncertainty is high (e.g., Jiang et al. (2023)). For repository-level code completion, these strategies are informative to some extent: in line completion and API completion from RepoEval, both heuristics can maintain the same level of performance with only 50% retrieval budget. However, we find that they fail to generalize well to all tasks and still incur a high latency cost as they need to conduct retrieval to make the decisions (Appendix C).
Instead, our framework adopts a self-selective RAG formulation. After observing $X_{l}$ and $X_{r}$, the LM directly self-triggers cross-file retrieval by generating a special token <cc> or abstains from retrieval via an empty token $\phi$ [^1] In practice, instead of greedily decoding <cc>, we check whether its probability exceeds a certain threshold.. This approach is inspired by the explorations in Kadavath et al. (2022), which show that an LM can be trained to predict whether it knows the answer or not without retrieval. Beyond this self-knowledge, our model also combines the question’s characteristics (i.e., whether retrieving cross-file information can likely help or not) in its judgment, as we will discuss in the next section. Finally, after the optional retrieval, the LM proceeds with the code completion with $X_{l}$, $X_{r}$, combined with $CC$ if retrieval is triggered.
Implementation-wise, self-selective RAG’s inference is conveniently modeled as an extension to fill-in-the-middle (Bavarian et al., 2022), with the entire process executed in a single left-to-right pass (Figure 2). One advantage of this design is the flexibility. The LM possesses the ability for RAG and fill-in-the-middle, and can seamlessly self-switch between the two when encountering different questions. Users can also easily adjust the ratio between the two through the retrieval threshold. Another advantage is its efficiency. The selective decision overhead is only a single forward pass, a significant save compared to making the retrieval decision via trial generation or trial retrieval. When the LM abstains from retrieval, it can directly proceed with generation and the retrieval overhead is completely avoided.
@@ 3.3 Self-supervised Multi-task Learning
    Retrieve $CC$ using the current file. We include $Y$ in the query for 50% of the data [^2] The main goal of the design is to align better with both non-iterative and iterative RAG use cases. During testing, a user may retrieve with both the in-file context and $Y’$ ’, a model’s draft prediction, which results in a $CC$ distribution close to that with $Y$ in the query (Zhang et al., 2023). .
@@
Figure 2: A comparison between fill-in-the-middle and self-selective RAG. We mark the end of the current file with a new token <eof>, which triggers the LM’s self-evaluation. $\rightarrow$ denotes the invocation of the LM. We color current-file context, retrieved contexts, and LM-generated parts in blue, green, and red respectively. fim_p, fim_s, and fim_m refer to the special tokens for fill-in-the-middle: fim_prefix, fim_suffix, and fim_middle. These tokens are already learned during the pre-training.
The full algorithms are presented in Appendix D. After running the algorithm, we obtain the fine-tuning instances, each in the form $(X_{l},X_{r},\ Y,\ CC,\ label)$.
@@ 4.1 Repoformer Implementation Details
We sample Python repositories from the Stack (Kocetkov et al., 2022). Basic filtering are applied to retain 18k repositories that have (1) at least five Python files, (2) at least three imports per file, and (3) at least two local imports per file. These criteria ensure the existence of local dependencies where RAG could be helpful. We use $\mathcal{M}$ = StarCoderBase-1B and $T$ = 00 to label 240k chunk and 120k function completion instances. We reserve 500 repositories for validation and use the rest for training.
@@
We fine-tune the 1B, 3B, 7B, and 16B variants of StarCoderBase with $\lambda=1.0$, maximum sequence length 2048, learning rate 2e-5, batch size 512, 50 warmup steps, and a linear learning rate decay. The models are trained for 2 epochs, which approximately takes 8, 12, 20, and 50 hours for the 1B/3B/7B/16B models respectively with 8 Nvidia A100 GPUs (40G memory). Our implementation is based on Jain et al. (2023) [^3] https://github.com/amazon-science/ContraCLM. We will call our models Repoformer-1B/3B/7B/16B. We have also applied the same method to train a multilingual version of Repoformer on a mixture of Python, Java, C#, and Typescript repositories. As we focus on the methodological discussion in the main text, we refer interested readers to Section E.2 for the detailed experiment setup and results.